OpenStudy (anonymous):

how do you factor 2x^2 + 9xy + 9y^2?

OpenStudy (cwrw238):

ill start you off add the first 2 equations to eliminate x ( 2x and -2x will disappear) y + z = 1.......................(1) now multiply equation 2 by -2 and add to equation 3 4x - 4y + 2z = 12 -4x + 3y - 2z = -9 add to eliminate x -y = 3 y = -3 now plug y = -3 in equation (1) above -3 + z = 1 z = 1+ 3 = 4 now plug y=-3 and z=4 into one of the original equations to find the value of x

OpenStudy (cwrw238):

you ok with that?

OpenStudy (anonymous):

yes thank you. could you help me with one more question? for some reason, this website won't let me ask a question right now. I have to solve for y and x for these equations, so I can graph them: 1. -3x + y = -6 2. y = -1/2 + 1

OpenStudy (cwrw238):

subtract to eliminate y: -3x = -6 +1/2 -1 -3x = - 6 1/2 = -13/2 x = 13/6 now plug y = 13/2 into equation 2 to find value of y